Circular No. 9021

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 D1 (LINEAR) J. V. Scotti, Lunar and Planetary Laboratory, University of Arizona,

reports his recovery of comet P/2001 CV_{8} (cf. IAUC 7581) on CCD images taken with the Spacewatch 1.8-m f/2.7 reflector. On the Feb. 19 images, Scotti noted the comet to be slightly diffuse with a 0′.3 tail in p.a. 295^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Feb. 19.53123 14^{h}25^{m}47.^{s}23 −18^{o}07′57′′.9 20.9 Scotti
19.54112 14 25 47.49 −18 08 01.9 20.8  _{”}
19.54978 14 25 47.71 −18 08 05.6 21.0  _{”}
20.49075 14 26 12.21 −18 14 49.9  _{”}
20.50000 14 26 12.30 −18 14 54.4 20.7  _{”}
20.50910 14 26 12.53 −18 14 59.1  _{”}

The indicated correction to the prediction on MPC 54170 is ΔT = −0.37 day (and to the prediction in the 2008/2009 Comet Handbook, ΔT = −0.24 day). The following orbital elements by B. G. Marsden are from 141 observations, 2001-2009 (mean residual 0′′.7):

Epoch = 2008 Oct. 21.0 TT

T = 2008 Oct. 10.9768 TT  ω = 151.^{o}5970
e = 0.444137  Ω = 359.8931 }2000.0
q = 2.159879 AU  i = 9.0351

a = 3.885631 AU n^{o} = 0.1286802 P = 7.66 years

COMETS C/2008 U5–U8 (SOHO) Further to IAUC 9020, additional near-sun presumed comets have been

found on SOHO website images — all Kreutz sungrazers except for C/2008 U6 (non-group), which was small and stellar in appearance (mag ∼ 7.5). C/2008 U5 was extremely faint (mag ∼ 8.5) and very diffuse. C/2008 U7 was stellar in appearance and perhaps slightly elongated (mag ∼ 7) in C3 images, but quite diffuse with a thin, diffuse, faint tail in C2 images. C/2008 U8 was very diffuse (mag ∼ 8–8.5).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 U5 Oct. 18.188 13^{h}28.^{m}1 −10^{o}55′ C2  RM 2009-B66
C/2008 U6 18.440 13 31.0 − 7 53 C2 MK 2009-B67
C/2008 U7  20.571 13 29.8 −12 09 C3/2 RK 2009-B67
C/2008 U8  20.746 13 36.0 −11 41 C2  TH 2009-B67
2009 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green